The pKa of NH4+ is 9.25 at 25°C. The rate constant at 25°C for the reaction of NH4+ and OH to form aqueous NH3 is 4.0 ✕ 1010 dm3 mol−1 s−1. Calculate the rate constant for proton transfer to NH3 (Units of s-1).
Answer #1

pKa ofNH4+

NH4+ + H2O <--> NH3 + H3O+

Ka = [NH3][H3O+]/[NH4+]

Ka = 10^-pKa = 10^-9.25

Let "K" be the rate constant for forward reaction and reverse(2nd order) reaction

k = 4*10^10 1/(Ms)

Calculte Rate constant for H+ transfer

K = [NH4+][OH-]/([NH3][H2O])

recall that

Kw = Ka*Kb

Kb = Kw/Ka = (10^-14)/(10^-9.25) = 1.78*10^-5

Then

for Rate constant

Kb = K/K'

K = Kb*K´

K = (1.78*10^-5)(4*10^10)

K = 7.12*10^5 1/s